OpenStudy (anonymous):

Compute the area of the following triangle. (Express answer in cm²) b = 20 in. h = 4 in. a = ____ cm2

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (mayankdevnani):

so ssimple

OpenStudy (mayankdevnani):

area of triangle=1/2*base*height

OpenStudy (mayankdevnani):

1/2*20*4=40cm^2=area

OpenStudy (mayankdevnani):

@Claire4christ ok

OpenStudy (anonymous):

ok so its 1/2x20x4

OpenStudy (mayankdevnani):

yaa

OpenStudy (mayankdevnani):

area=40cm^2

OpenStudy (anonymous):

1,600

OpenStudy (mayankdevnani):

what?

OpenStudy (anonymous):

so its ss40x40 basically

OpenStudy (anonymous):

The initial units are in inches and the answer needs to be converted to cm. Or convert the initial units to cm before using 1/2 x b x h

OpenStudy (anonymous):

how you do that i don'n remember

OpenStudy (anonymous):

Lets convert to cm first. Do you know how many cm in an inch? One of us can look it up.

OpenStudy (anonymous):

50.8=20in

OpenStudy (anonymous):

1 inch = 2.54 cm

OpenStudy (anonymous):

20 inch = 20 x 2.54 cm

OpenStudy (anonymous):

b=50.8

OpenStudy (anonymous):

OK so just use the 1/2 x b x h with each measure converted to cm and "ta da."

OpenStudy (anonymous):

518.16

OpenStudy (anonymous):

259.08

OpenStudy (anonymous):

cm2

OpenStudy (anonymous):

259.08x259.08

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Wait, wait. The units are squared. you don't square you answer. I'll try to explain in a second.

OpenStudy (anonymous):

so its gonna be 259.08

OpenStudy (anonymous):

259

OpenStudy (anonymous):



OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Smae for a triangle or any shape. You use a formula or equation to figure out the area and the units of an area are squares, literally, a 4 sided square, 1 unit by 1 unit, 1 square unit.

OpenStudy (anonymous):

Do you understand? This is pretty important to know.

OpenStudy (anonymous):

ok so we convert them

OpenStudy (anonymous):

and i got 259.08

OpenStudy (anonymous):

259.08 what?

OpenStudy (anonymous):

inches, centimeters, square inches or square centimeters?

OpenStudy (anonymous):

cm2

OpenStudy (anonymous):

What I want to make sure you understand is what the area is and the units of area measurement, what they are and how you get them. If you got it, great.

OpenStudy (anonymous):

yes i got that

OpenStudy (anonymous):

is my answer right or not

OpenStudy (anonymous):

squaring your answer after calculating the area is what concerned me.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

you told me not do it.

OpenStudy (anonymous):

its says not right

OpenStudy (anonymous):

20x2.5=50 4x2.5=10

OpenStudy (anonymous):

b = 20 in = 20 x 2.54 cm = 50.8 cm h = 4 in = 4 x 2.54 cm = 10.16 cm Area = 1/2 x 50.8 cm x 10.16 cm Area = 1/2 (516.128) cm^2 = 258.064 cm^2 Which is completely precise. Do they tell you to round?

OpenStudy (anonymous):

i got 250

OpenStudy (anonymous):

If they are using 2.5 cm per inch, that would make it less precise. But if that is what they used then that is what you should use. Is 250 cm^2 correct?

OpenStudy (anonymous):

not right weird

OpenStudy (anonymous):

I am clueless. Sorry. You are doing it right. 250 is what I get if 1 in = 2.5 cm.

OpenStudy (anonymous):

ill ask my teacher about that question tomorrow

OpenStudy (anonymous):

If rounding to one decimal place were to happen after calculating using 1 in = 2.54 cm then the answer would be 258.1 cm^2 not 259.08. 259.08 is a little bit off, maybe that's it.

OpenStudy (anonymous):

still not right that's ok ill ask my teacher tommorow.